Solve the equation 2X+4y=10and3X+6Y=15
Answer by stanbon(75887)

Solve the system of equations:
2X+4y=10
3X+6Y=15
-----
Multiply thru the 1st by 3 ; Multiply thru the 2nd by 2:::
6x + 12y = 30
6x + 12y = 30
========
The two equations are really the same.
Solution: All points (x ,(-1/2)x+(5/2)) where x is any Real Number.
================
Cheers,
Stan H.